Find the three current I1, I2, I3 with Kirchoff rule where R1= 100ohms, R2= 10ohms, R3= 5ohms and E1= 12volts, E2= 6volts


1
Expert's answer
2022-02-24T11:33:02-0500

Answer

Using ohm's law

Curents

"I_1=\frac{E_1}{R_1}=\frac{12}{100}=0.12A"

"I_2=\frac{E_2}{R_2}=\frac{6}{10}=0.6A"

So

"I_3=I_1+I_2=0.12+0.6=0.18A"